[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Oct 17, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Oct 18 04:14:57 UTC 2017


HSU Shift Log for Wed Oct 18 2017 00:14:48 GMT-0400 (EDT)

Shift: R

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: GOES-West anomaly/outage was reported from 2259Z to  
0052Z (no images received between those times).

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171018/69bcbba0/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list